OpenStudy (anonymous):

Evaluate: 32 ÷ (2 + 1)

OpenStudy (anonymous):

10.66666.........

OpenStudy (anonymous):

Which of the following pairs of terms are like terms? 2x and 3y 2x and 4x2 x and x2 5x3 and 6x3

OpenStudy (anonymous):

\[\frac{32}{2+1}=\frac{32}{3}=10+\frac{2}{3} \]